1. Which of the following is a weak acid? (a) HNO3 (b) HI (c) HBr (d) HF (e) HClO3 2. Which salt is not derived from a strong acid and a strong soluble base? (a) MgCl2 (b) Ba(NO3)2 (c) LiClO4 (d) CsBr (e) Nal. 3. Which of the following is a strong electrolyte? (a) H2O (b) KF (c) HF (d) HNO2 (e) (CH3)3N. 4. In a sample of pure water, only one statements is always true at all conditions of temperature and pressure. Which one is always true? (a) [H3O^+] = 1.0x10^-7 M (b) [OH^-] = 1.0x10^-7 M (c) PH = 7.0 (d) pOH = 7.0 (e) [H3O^+] = [OH^-] 5. If Kw is 2.9x10^-15 at 10 degrees Celsius, what is the pH of water at 10 degrees Celsius? 6. The pOH of a solution of NaOH is 11.30. What is the [H^+] for this solution? 7. The [H3O^+] in a 0.050 M solution of Ba(OH)2 is: 8. What is the approximate pH of a solution labeled 6x10^-5 M HBr?
1
Expert's answer
2020-06-16T15:03:24-0400
  1. d
  2. c
  3. b
  4. e
  5. 7 since pH=pOH in water
  6. 2,7 because pH=14-pOH. [H+]= 10-2,7=0,00199
  7. Ba(OH)2"\to" Ba+2 + 2OH- =>[OH-]=0,01=> pOH=2=>pH=12
  8. HBr"\to" H++Br- pH=-lg[H+]=4,22